OpenStudy (anonymous):

solve with a matrix equation : 4x-3y=15 2x+y=5

OpenStudy (fibonaccichick666):

so, can you set up the matrix of coefficients?

OpenStudy (anonymous):

[4 -3] [2 1]

OpenStudy (anonymous):

[4 -3][x]= [15] [2 1] [y] [5]

OpenStudy (michele_laino):

please solve the second equation for y

OpenStudy (michele_laino):

y=5-..., please complete

OpenStudy (anonymous):

y=5-2x

OpenStudy (michele_laino):

ok!, now insert your y-value into the first equation

OpenStudy (loser66):

@Michele_Laino but the question is about "solve with a matrix"

OpenStudy (anonymous):

4x-3(5-2x)=15

OpenStudy (anonymous):

x=3

OpenStudy (michele_laino):

@Loser66 you are right!

OpenStudy (fibonaccichick666):

sorry, was at another spot, so the first thing you gave me: was nearly correct "[4 -3] [2 1] "

OpenStudy (fibonaccichick666):

you just left out the right side of the eq :)

OpenStudy (fibonaccichick666):

can you try to rewrite it now?

OpenStudy (michele_laino):

@adjsr1234 that's right!, now insert x=3, into your y-value

OpenStudy (michele_laino):

in order to solve your system using matrices, please note taht your system is:

OpenStudy (anonymous):

My answer choices are not in matrix form so I figured out the answer. I got : (3, -1)

OpenStudy (fibonaccichick666):

but you should learn how to do the problem

OpenStudy (michele_laino):